# Reducing Vision Transformer Latency on Edge Devices via GPU Tail Effect and Training-free Token Pruning

### v1

@@ 3.3 Token Pruning Method
In terms of design goals, we want our pruning mechanism to require (1) no training or finetuning on $M$, and (2) to introduce minimal overhead. First, requiring no fine-tuning or training is a convenience for practitioners, and allows pre-trained models to be deployed readily. Second, keeping overhead to a minimum for small workload sizes is necessary to yield latency reductions as previously mentioned in 3.2.
The primary design choice of pruning techniques is choosing a method to rank the importance of tokens. Similarly to other pruning techniques [3, 14], we rank the importance of each token $i$ by measuring the quantity of attention received from all other tokens. This is analogous to summing along the columns of the attention matrix in a transformer encoder layer.
Our method incorporates two aspects of importance, one from the $A_{\text{H}}$ matrix and one from the $V$ matrix. For attention, we proceed as follows. Recall that the core attention mechanism of transformers [35] is parameterized by the $Q$, $K$, and $V$ matrices. The attention matrix (including softmax) of a transformer encoder layer with $N$ tokens and $H$ heads is given by $A_{\text{H}}\in\mathbb{R}^{H\times N\times N}$  [35]. To convert the softmax attention matrix, $A_{\text{H}}$, into tokenwise importance scores we apply a reduction along the along the head dimension, $H$. We find the maximum along the head dimension for $A_{\text{H}}$. We found that applying max marginally increased accuracy compared to mean or sum. This reduction yields a new matrix $A\in\mathbb{R}^{N\times N}$. Now, summing across rows of $A$ yields the quantity of attention each token $i$ receives, similar to a component of EViT’s pruning mechanism [16] or Top-K [12]. This yields an importance score $\text{A}_{\text{S}}[i]\in\mathbb{R}^{N}$ for each token $i$.
$V_{\text{S}}$ Feature Choice Feature ImageNet1K Accuracy (%) Q 79.74 K 79.77 V 79.81
$V_{\text{S}}$ Head Reduction Operator Feature ImageNet1K Accuracy (%) mean 79.72 sum 79.70 max 79.81
Table 2: Token pruning design ablations using for our $V_{S}$ term our schedule on DeiT-S-dist for batch-size 2. In our code we use torch.amax [29] instead of max to handle the batch dimension.
@@
We then consider a second importance term based on the $V\in\mathbb{R}^{H\times N\times C}$ matrix. This term assigns higher scores to tokens whose entry in the $V$ matrix has a large magnitude compared with other tokens. Adding this term can marginally improve accuracy. Similarly to the attention matrix, we find the maximum along the $H$ dimension [29] — we do a set of ablations summarized in Tab. 2. Post reduction, we have $V^{\prime}\in\mathbb{R}^{N\times C}$ and we have:
$\text{V}_{\text{S}}[i]$ $=\text{softmax}(\sum^{C}_{c=1}V^{\prime}[i,c])\in\mathbb{R}^{N}.$ (5)
We add both of these scores to yield our importance score for each token $i$:
$\text{Importance}[i]$ $=\text{A}_{\text{S}}[i]+\text{V}_{\text{S}}[i].$ (6)
Finally, we borrow from EViT [16] and instead of discarding the pruned tokens, we create a new “inattentive” token based on the averaged features of all pruned tokens and append it to the set of kept tokens. This allows information from pruned tokens to be preserved in this inattentive token, which EViT found improved accuracy. Fig. 5 illustrates the token pruning process for a ViT encoder block or layer. Note that we do not prune specialized tokens such as the CLS [9] or DIST [33] tokens. These tokens serve additional purposes beyond representing the input.
@@ 4 Evaluation
In this section, we demonstrate the effectiveness of our token pruning schedule and pruning mechanism across devices (listed in Tab. 3) and across various workload sizes. To justify our design decisions, we investigate our $\alpha$ hyperparameter and pruning location choice for our schedule selection and pruning mechanism in Sec. 4.2 and Sec. 4.3, respectively. Then, we compare with a variety of existing token removal methods such as DynamicViT [30], ToMe [1]. We use Top-K [16, 12] as a baseline on ImageNet1K [8]. Sec. 4.4 demonstrates how our method advances the accuracy vs. latency pareto front across devices, workload sizes, and models. Finally, we identify some pitfalls of existing token pruning techniques for small workload sizes in Sec. 4.5. Particularly we show that some existing methods actually increase latency rather than reduce it.
@@ 4.1 Experimental Setup
Evaluation is done using three devices with varying hardware characteristics (Tab. 3). These devices feature various generations of NVIDIA GPU architectures, and thus feature different levels of computing power. We use NVIDIA TX2 [24] and NVIDIA AGX Orin [21] development boards which are edge devices designed for machine learning. Additionally, evaluation is done on a server-grade system with an NVIDIA A100 GPU [22] to demonstrate our method can generalize across a wide range of hardware, even though we design our method for edge devices. Experiments performed on the TX2 and AGX Orin were performed with fixed CPU and GPU clock rates for reproducibility and consistency. The A100 system clocks were unable to be locked because the servers machines are university-wide resources.
Device RAM (GB) GPU Cores CUDA Cores Max Power (W)
TX2 [24] 8 2 256 15
AGX Orin [21] 32 14 1792 40
A100 [22]  ✝ 80 108 6192 300
In order to measure latency for evaluation, we use the PyTorch benchmark module [28]. All accuracy measurements were performed on the A100 system using a batch size of 512. Conveniently, accuracy only needs to be measured once for a particular pre-trained model $M$. It is only necessary to record the latency $L(n)$ on a target device for a particular workload size in order to capture GPU tail effect behavior. In subsequent experiments, we decide the pruning hyperparameters of each token pruning method for fair comparison with our method.
For reproducibility, our code is open source. We include our method, and modifications to the original ToMe codebase in order to support DinoV2 pre-trained models.
@@ 4.2 Pruning schedule ablation study across $\alpha$ (utility weight of latency/accuracy)
In 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect. The hyperparameter of this algorithm is $\alpha$, which determines the relative weighting of the accuracy and latency utilities. Setting to $\alpha=1$ prioritizes accuracy, and only a few tokens might be pruned. Setting to $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens. To understand the performance spectrum, we perform an ablation study. The data is in Tab. 4 and illustrated in Fig. 6.
Device $\alpha$ Range Computed $R$ Accuracy Median
(# tok. to remove) Degradation Latency (ms)
AGX Orin 0.1-0.3 139 1.07 112.2 (-27.8%)
0.4-0.5 166 2.04 104.2 (-33.0%)
0.6-0.9 193 4.82 100.9 (-35.1%)
A100 0.1-0.4 73 0.25 28.77 (-4.70%)
0.5-0.9 139 1.07 28.27 (-6.34%)
Table 4: A companion table for Fig. 6 that shows the accuracy vs latency tradeoffs for the computed number of tokens to prune $R$ for various $\alpha$. Latency percent change is computed with respect to the unmodified DinoV2-G inference time.
Figure 6: Computed pruning rate $R$ from our offline pruning schedule versus various $\alpha$. Note that the selection of $R$ is relatively invariant to changes in $\alpha$. Configuration: $M$ = DinoV2-G and batch size = 2.
@@ 4.3 Pruning location ablation study
Batch Size Pruning Accuracy Median
Layer Degradation Latency (ms)
1 1/40 3.19 68.4
@@ 4.4 Accuracy-latency tradeoffs of our method
First, our method is able to achieve higher accuracy than other training-free pruning techniques ToMe and Top-K. Tab. 6 is an ablation study across various workload sizes (batch size, models) and devices, where we tune the hyperparameters of methods such that similar latency is achieved. Unsurprisingly, DynamicViT is able to retain higher accuracy than all methods, since it requires training.
Device Batch Model and Accuracy Median
Size Method Degradation Latency (ms)
TX2 2 DeiT-B + Top-K 2.44 143.0 (-33.5%)
DeiT-B + ToMe 1.63 142.7 (-33.7%)
DeiT-B + Ours 1.16 146.0 (-32.1%)
DeiT-B + DyViT ✝ 0.57 143.7 (-33.2%)
DeiT-B 215.0 -(00.0%)
TX2 2 DeiT-S + Top-K 4.30 50.32 (-27.0%)
DeiT-S + ToMe 2.47 52.98 (-23.1%)
DeiT-S + Ours 1.24 49.44 (-28.3%)
DeiT-S + DyViT ✝ 0.46 50.88 (-26.2%)
DeiT-S 68.92 -(00.0%)
AGX Orin 2 DinoV2-G + Top-K 45.66 104.4 (-32.9%)
DinoV2-G + ToMe 6.96 104.4 (-32.9%)
DinoV2-G + Ours 2.04 104.1 (-33.1%)
DinoV2-G 155.5 -(00.0%)
A100 4 DinoV2-G + Top-K 13.31 33.76 (-16.7%)
DinoV2-G + ToMe 6.96 33.79 (-16.6%)
DinoV2-G + Ours 2.04 32.37 (-20.1%)
DinoV2-G 40.53 -(00.0%)
Compared with ToMe and Top-K however, it is shown that our method yields consistently lower accuracy degradation. Batch size 2 inference on the AGX Orin demonstrates that to achieve similar latency to our method, Top-K degrades accuracy by more than 45.6%, while we degrade accuracy by only 2%. Across all workload sizes and devices, ToMe featured 0.47%-5.92% lower accuracy than our method (Tab. 6).
Figure 7: Illustration of accuracy and latency tradeoffs of surveyed methods with $M$ = DinoV2-G. Our pruning schedule selection pruning mechanism generate points that expand the pareto front. This indicates that considering the GPU tail effect can provide strictly better tradeoffs than existing methods in some cases. Note that the number of tokens removed at each layer ($r$) of Top-K and ToMe is evaluated from $r=5$ to $=8$ in increments of 1.
Second, we note that for larger pruning rates such as $r=7$ and $r=8$, ToMe and Top-K remove nearly all input tokens by the last layer of DinoV2-G. These high pruning ratios yield high accuracy degradation of over 40% in the case of A100 batch-size 4 for ToMe, as seen in Fig. 7. Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we simultaneously achieve higher accuracy and lower latency than progressive pruning approaches. This shows aggressive pruning at early layers can be strictly better than progressive pruning methods for certain workload sizes. Our token pruning mechanism and schedule offers more accuracy latency tradeoffs across workload sizes.
@@ 4.5 Low workload size observations
Device Batch Model and Accuracy Median
Size Method Degradation Latency (ms)
TX2 1 DeiT-S + Top-K (r=6) 0.73 43.38 (+18.6%)
DeiT-S + ToMe (r=6) 0.27 50.70 (+30.3%)
DeiT-S + Ours (R=69) 0.99 32.12 (-9.06%)
DeiT-S 35.32 (-0.00%)
Table 7: Given comparable total pruned tokens (69 to 72), for certain workload sizes a token pruning mechanism may increase latency due to the overhead of performing pruning at multiple layers. Our pruning method is applied at a single layer to avoid this relatively significant overhead.
@@ 5 Conclusion
In this work, we investigated how to improve the efficiency of pre-trained vision transformers when deployed on edge devices and provided three primary contributions. First, we provided insight into how to determine a token pruning schedule for transformer models by leveraging information about the GPU tail effect. Second, a token pruning approach was introduced that is training-free and has low overhead compared with other pruning methods. We show that our method advances the accuracy-latency pareto front compared with existing token pruning approaches across workload sizes and devices. Furthermore we show that for single batch inference, other methods may actually increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%. For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
6 Limitations
Our approach to token pruning is designed to achieve better tradeoffs at small workload sizes than other methods. This is ideal when inputs are streamed to the model, or if only small batch inference is possible. For larger workload sizes, such as running DinoV2-G with batch sizes of 32 or higher, our method may not be ideal.
7 Future Work
We demonstrate that single layer pruning can be effective even for earlier pruning layers. Learning based approaches may be able to take advantage of the high latency reductions of aggressive, early pruning while maintaining accuracy. Another promising direction is investigating the benefits of composing token sparsification and other acceleration methods together.

### v2

@@ 3.3 Token Pruning Method
In terms of design goals, we want our pruning mechanism to require (1) no finetuning on $M$ and (2) to introduce minimal overhead. First, requiring no fine-tuning or training is a convenience for practitioners, and allows pre-trained models to be deployed readily. Second, keeping overhead to a minimum for small workload sizes is necessary to yield latency reductions as previously mentioned in 3.2.
The primary design choice of pruning techniques is choosing a method to rank the importance of tokens. Similarly to other pruning techniques [3, 15], we rank the importance of each token by measuring the quantity of attention received from all other tokens.
@@
Our method incorporates two aspects of importance, one from the $A_{\text{H}}$ matrix and one from the $V$ matrix. For the $A_{\text{H}}$ term, we proceed as follows. Recall that the core attention mechanism of transformers [40] is parameterized by the $Q$, $K$, and $V$ matrices. The attention matrix (including softmax) of a transformer encoder layer with $N$ tokens and $H$ heads is given by $A_{\text{H}}\in\mathbb{R}^{H\times N\times N}$  [40]. To convert the softmax attention matrix $A_{\text{H}}$ into tokenwise scores we: (1) reduce the head dimension $H$ with amax [35] (2) sum along the row $N$ dimension (3) normalize scores by dividing by the max. This transforms $A_{\text{H}}$ into tokenwise scores $\text{A}_{\text{S}}\in\mathbb{R}^{N}$, similar to a component of EViT’s pruning mechanism [19] or Top-K [13].
We then consider a second importance term based on the $V\in\mathbb{R}^{H\times N\times C}$ matrix. This term assigns higher scores to tokens who have high weighting in the $V$ matrix, which marginally increased accuracy. We perform a similar operation for this term: (1) reduce the $H$ dimension of $V$ with amax [35] (2) sum along the feature dimension $C$ (3) normalize scores with softmax. The resulting score vector is $V_{\text{S}}\in\mathbb{R}^{N}$, which we add to $\text{A}_{\text{S}}$ for our overall score:
Token Importance $=\text{A}_{\text{S}}+\text{V}_{\text{S}}.$ (4)
Finally, we borrow from EViT [19] and instead of discarding the pruned tokens, we create a new “inattentive” token based on the averaged features of all pruned tokens and append it to the set of kept tokens. This allows information from pruned tokens to be preserved in this inattentive token, which EViT found improved accuracy. Fig. 5 illustrates the token pruning process for a ViT encoder block or layer.
@@ 4 Evaluation
In this section, we demonstrate the effectiveness of our token pruning schedule and pruning mechanism across devices (listed in Tab. 3) and across various workload sizes. We primarily focus on smaller workloads, since these are representative of edge deployment use-cases [18]. To justify our design decisions, we investigate our $\alpha$ hyperparameter and pruning location choice for our schedule selection and pruning mechanism in Sec. 4.2 and Sec. 4.3, respectively. Then, we compare with a variety of existing token removal methods such as DynamicViT [36], ToMe [1]. We use Top-K [19, 13] as a baseline on ImageNet1K [9]. Sec. 4.4 demonstrates how our method advances the accuracy vs. latency pareto front across devices, workload sizes, and models. Finally, we identify some pitfalls of existing token pruning techniques for small workload sizes in Sec. 4.5. Particularly we show that some existing methods actually increase latency rather than reduce it.
Device GPU CUDA Max Power
Cores Cores (W)
TX2 [29] 2 256 15
Orin [24] 14 1792 40
A100 [25]  ✝ 108 6192 300
@@ 4.1 Experimental Setup
Evaluation is done using three devices with varying hardware characteristics (Tab. 3). We use NVIDIA TX2 [29] and NVIDIA AGX Orin [24] development boards which are edge devices designed for machine learning. Additionally, evaluation is done on a server-grade system with an NVIDIA A100 GPU [25] to demonstrate our method can generalize across a wide range of hardware, even though we design our method for edge devices. Experiments performed on the TX2 and Orin were performed with fixed CPU and GPU clock rates for reproducibility and consistency. The A100 system clocks were unable to be locked because the servers machines are university-wide resources.
In order to measure latency for evaluation, we use the PyTorch benchmark module [33]. All accuracy measurements were performed on the A100 system with batch size 512 on the evaluation split of ImageNet1K. In subsequent experiments, we decide the pruning hyperparameters of each token pruning method for fair comparison with our method. For reproducibility, our code is open source.
@@ 4.2 Pruning schedule ablation study across $\alpha$ (utility weight of latency/accuracy)
In 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect.
Device $\alpha$ R Acc. Median
Range Loss Latency (ms)
Orin [0.1, 0.3]0.10.3[0.1,\;0.3][ 0.1 , 0.3 ] 139 1.07 (-27.8%) 112.2
[0.4, 0.5]0.40.5[0.4,\;0.5][ 0.4 , 0.5 ] 166 2.04 (-33.0%) 104.2
[0.6, 0.9]0.60.9[0.6,\;0.9][ 0.6 , 0.9 ] 193 4.82 (-35.1%) 100.9
A100 [0.1, 0.4]0.10.4[0.1,\;0.4][ 0.1 , 0.4 ] 73 0.25 (-4.70%) 28.77
[0.5, 0.9]0.50.9[0.5,\;0.9][ 0.5 , 0.9 ] 139 1.07 (-6.34%) 28.27
Table 4: Accuracy vs latency tradeoffs for the computed number of tokens to prune $R$ for various $\alpha$ of our method. Latency percent change is with respect to baseline DinoV2-G inference time.
The hyperparameter of this algorithm is $\alpha$, which determines the relative weighting of the accuracy and latency utilities. Setting to $\alpha=1$ prioritizes accuracy, and only a few tokens might be pruned. Setting to $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens. To understand the performance spectrum, we performed an ablation study. The data is in Tab. 4.
@@ 4.3 Pruning location ablation study
Batch Pruning Acc. Median
Size Layer Loss Latency (ms)
1     01/40 3.19 68.4
@@
Device Batch Model & Acc. Median
Size Method Loss Latency (ms)
TX2 2 DeiT-B -(00.0%) 215.0
w/ Top-K 2.44 (-33.5%) 143.0
w/ ToMe 1.63 (-33.7%) 142.7
w/ DyViT ✝ 0.57 (-33.2%) 143.7
w/ Ours 1.16 (-32.1%) 146.0
TX2 2 DeiT-S -(00.0%) 68.92
w/ Top-K 4.30 (-27.0%) 50.32
w/ ToMe 2.47 (-23.1%) 52.98
w/ DyViT ✝ 0.46 (-26.2%) 50.88
w/ Ours 1.24 (-28.3%) 49.44
Orin 2 DinoV2-G -(00.0%) 155.5
w/ Top-K 45.66 (-32.9%) 104.4
w/ ToMe 6.96 (-32.9%) 104.4
w/ Ours 2.04 (-33.1%) 104.1
A100 4 DinoV2-G -(00.0%) 40.53
w/ Top-K 13.31 (-16.7%) 33.76
w/ ToMe 6.96 (-16.6%) 33.79
w/ Ours 2.04 (-20.1%) 32.37
@@ 4.4 Accuracy-latency tradeoffs of our method
First, our method is able to achieve higher accuracy than other training-free pruning techniques ToMe and Top-K. Tab. 6 is an ablation study across various workload sizes (batch size, models) and devices, where we tune the hyperparameters of methods such that similar latency is achieved. Unsurprisingly, DynamicViT is able to retain higher accuracy than all methods, since it requires training. Compared with ToMe and Top-K however, it is shown that our method yields consistently lower accuracy degradation. Batch size 2 inference on the AGX Orin demonstrates that to achieve similar latency to our method, Top-K degrades accuracy by more than 45.6%, while we degrade accuracy by only 2%. Across all workload sizes and devices, ToMe featured 0.47%-5.92% lower accuracy than our method (Tab. 6).
Figure 6: Illustration of accuracy-latency tradeoffs of surveyed methods with $M$ = DinoV2-G: (a) batch size=2 on AGX Orin (b) batch size=4 on A100 (c) batch size=4 on AGX Orin. Our pruning schedule and mechanism generate points that expand the pareto front. The number of tokens removed at each layer ($r$) of Top-K and ToMe is evaluated from $r=5$ to $=8$ in increments of 1.
Second, we note that for larger pruning rates such as $r=7$ and $r=8$, ToMe and Top-K remove nearly all input tokens by the last layer of DinoV2-G. These high pruning ratios yield high accuracy degradation of over 40% in the case of A100 batch-size 4 for ToMe, as seen in Fig. 6. Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we achieve higher accuracy and lower latency than progressive pruning approaches. This shows aggressive pruning at early layers can be strictly better than progressive pruning methods for certain workload sizes.
@@ 4.5 Low workload size observations
Device Method Acc. Median
Loss Latency (ms)
TX2 Baseline (-0.00%) 35.32
Top-K (r=6) 0.73 (+18.6%) 43.38
ToMe (r=6) 0.27 (+30.3%) 50.70
Ours (R=69) 0.99 (-9.06%) 32.12
Table 7: Experiment on DeiT-S with batch size 1 that illustrates pruning overhead. Given comparable total pruned tokens (69 to 72), for certain workload sizes a token pruning mechanism may increase latency due to the overhead of performing pruning at multiple layers. Our pruning method is applied at a single layer to avoid this relatively significant overhead.
@@
5 Limitations
Our approach to token pruning is designed to achieve better tradeoffs when latency-workload is non-linear. This occurs primarily for small workload sizes and batch sizes, which are common contexts in edge deployment. For large workloads such as DinoV2-G with batch size 256+, our method is less effective because latency-workload relationships becomes linear. Last, though we avoid training, our method requires a grid-search. In some contexts, this may be expensive (many tokens, large batch size, large model). These contexts are outside the focus of this work, but may be relevant for some practitioners.
@@ 6 Conclusion
In this work, we investigated how to improve the efficiency of vision transformers by considering latency-workload relationships. We made three main contributions. First, we provide insight on how ViT latency is affected by underlying hardware, machine learning framework, and kernel launch factors. Second, we show how to determine a token pruning schedule by leveraging information about latency-workload relationships. Third, we introduced a token pruning mechanism that is training-free and low overhead. Our method advances the accuracy-latency Pareto front compared with existing token pruning approaches across workload sizes and devices. For single batch inference, other methods may increase latency by 18.6-30.3% with respect to baseline, while we reduce it by 9%. For similar latency (within 5.2%) across devices, the state-of-the-art, ToMe, achieves 45.8%-85.4% ImageNet1K accuracy, while we achieve 78.6%-84.5%. Improving kernel designs to reduce non-linear latency behavior is a promising future direction.